Maggie is going to the mall with her friends and has $200 to spend from her birthday and Christmas money. She finds a store that has jeans for $25 and dresses for $50 (prices include tax).
When Maggie leaves the store she has spent all of her $200 and has bought 6 items of clothing. How many pairs of jeans and many dresses did she purchase?

Answer by TimothyLamb(4379)

x = pairs of jeans
y = number of dresses
---
x + y = 6
25x + 50y = 200
---
put the system of linear equations into standard form
---
x + y = 6
25x + 50y = 200
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 4
y = 2
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---